Which of the following expressions represents the correct expansion of the y determinant for the equations:
x+y=0
x+z=4
y+z=0

Answer by TimothyLamb(4379)

x+y=0
x+z=4
y+z=0
---
put the system of linear equations into standard form
---
x+y=0
x+z=4
y+z=0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 2
y = -2
z = 2
---
system matrix
1 1 0
1 0 1
0 1 1
---
inverse of system matrix
0.5 0.5 -0.5
0.5 -0.5 0.5
-0.5 0.5 0.5
---
determinant of system matrix = -2
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---